# A Collectivist, Economic Perspective on AI

### v1

@@ 1 Introduction
The study of multiple agents in computer science is by no means new, and researchers in fields such as multi-agent ML, human-computer interaction, and algorithmic game theory will recognize that it is their work that is being promoted here, as are the perspectives of the social sciences broadly speaking. Moreover, there are antecedents of our arguments in the study of “collective intelligence” (Tumer and Wolpert, 2004; Parkes, 2004; Malone and Bernstein, 2015). That literature has two main branches, one focused on qualitative, experimental work with groups, and the other focused on artificial collectives with designed agent utilities. Our focus is different. We want to uncover quantitative design principles for emerging real-world ML-based systems in which many of the participants are human and many are non-human. The goals and utilities of the humans are to be understood and respected, not designed. The principles should be simultaneously computational, economic, and inferential.
Of particular importance, we view economics both as a source of distributed algorithms (“mechanisms”) and also as a social science. We argue that to build AI systems that enhance the human condition, while operating in our social world, we need to make full use of economic ideas such as markets, incentives, information design, trade, prices, contracts, and externalities. In particular, markets, which are collectives that are based on certain patterns of interactions and information flows, can be creative, robust, and scalable. They can operate over long spans of time, and they do more than just generate text and images—they make real-world decisions, create value, spread value among multiple participants, and hedge against risk. They are, in a word, intelligent. Indeed, they were intelligent even before the advent of computing, and it seems likely that they can become even more intelligent—in hitherto unexplored ways—if we view economics through the lens of modern ML (and vice versa). Moreover, social intelligence can be naturally aligned with human welfare.
In summary, our goal is to connect economic ideas—in particular, the fields of mechanism design (Hurwicz and Reiter, 2006) and information design (Bergemann and Morris, 2019)—with the data-analytic and large-scale perspectives of modern ML. We want designers to speak not only the language of learning algorithms but also be able to consider the overall social system in which ML algorithms are embedded and to treat social welfare as a first-class citizen in their design. Developing a collectivist perspective on information technology that connects algorithmic and statistical concepts to economic and social concepts can be just as exciting intellectually as AGI, and at least as promising for the future of the species.
2 Uncertainty, Collectives, and LLMs
LLMs can give the impression that they are quantifying uncertainty and reasoning inductively. But this is in part an illusion, arising from the fact that the humans who produced the data on which the learning algorithms are based employed terminology and arguments associated with reasoning under uncertainty. Real-world management of uncertainty needs a more solid foundation.
Uncertainty is a broad concept, encompassing the kinds of sampling-based and probabilistic assertions one finds in statistics, the epistemic considerations of the sciences, and the information asymmetries of economics. Moreover, real-world uncertainty is shaped by phenomena such as the provenance of data and the freshness of data. A doctor relying on medical data that is several years old should be more uncertain about a conclusion than someone relying on fresh data, all else being equal. Such judgments should trade off with respect to other sources of uncertainty. LLMs are currently far from humans with respect to this kind of real-world uncertainty management. Even once known technical issues are mitigated in LLMs, such as systematic overconfidence (Sun et al., 2025), and even once basic statistical concepts are more fully incorporated in LLM design (Ji et al., 2025), LLMs will continue to fail to be trustworthy for high-stakes decision-making until the management of uncertainty becomes a first-class citizen in the LLM research agenda.
@@ 2 Uncertainty, Collectives, and LLMs
A broad understanding of uncertainty in human intelligence goes beyond cognitive science into social science. Indeed, collectives play an essential role in changing the world in ways that shape and mitigate uncertainty. An individual human foraging for fruit may or may not succeed in finding fruit on any given day, but once a collective creates a market for produce, the uncertainty regarding finding fruit drops significantly. A human can then depend on fruit being available and build on that certainty, perhaps opening a pastry shop.

### v2

@@ 5 Discussion
The study of multiple agents in computer science is by no means new, and researchers in fields such as multi-agent ML, human-computer interaction, and algorithmic game theory will recognize that it is their work that is being promoted here, as are the perspectives of the social sciences broadly speaking. Moreover, there are antecedents of our arguments in the study of “collective intelligence” (Tumer and Wolpert, 2004; Parkes, 2004; Malone and Bernstein, 2015). That literature has two main branches, one focused on qualitative, experimental work with groups, and the other focused on artificial collectives with designed agent utilities. My focus is different. I want to uncover algorithmic design principles for emerging real-world ML-based systems in which many of the participants are human and many are non-human. The goals and utilities of the humans are to be understood and respected, not designed. The principles should be simultaneously computational, economic, and inferential.
Although ML-powered markets can be expected to differ in important ways from classical markets, it’s worthwhile to recall some of the appealing features of classical markets. First, they provide a form of uncertainty reduction—buyers can count on products to be available and plan accordingly. Second, they cope with heterogeneity. Third, they create new roles on an as-needed basis. It is early days for ML-powered markets, but it is not hard to imagine many new roles that can be expected to arise, akin to those that have arisen in previous eras of technology but based around data and learning—auditors, brokers, aggregators, sellers, buyers, artists, forecasters, insurers, and explorers. These roles will give rise to personalized services, economies of scale, and appropriate touch points at which regulatory control can be exerted to mitigate those biases that are proscribed by legal or ethical considerations.
Design principles based on our tripartite blend make it easier to discuss issues of interest to both individuals and collectives, such as privacy, fairness, ownership, alignment, reputation, and transparency. The blend allows these issues to be treated as tradeoffs rather than being reduced to black-and-white distinctions. For example, differential privacy is a computational method that adds noise to data to guarantee privacy according to a particular definition (Dwork and Roth, 2014). But this is only part of the story. An individual’s decision to ask for a certain level of differentiable privacy in an interaction will trade off with the other costs and benefits that the individual expects from the interaction. Moreover, inferential issues arise whenever noise is being added to data, and there are quantitative tradeoffs that can be formulated that link inferential accuracy and privacy (Duchi et al., 2014). Making full use of the tripartite blend will generally involve solutions that take the form of tradeoffs.
Finally, for AI to grow into a mature engineering discipline that delivers systems that yield value to humans in real-world settings, it will need far more than blends of existing algorithms (and far more than just more data and more compute). In this regard, it’s useful to learn from the history of the fields of chemical engineering and electrical engineering, which brought the complex phenomena of chemical reactions and electromagnetism under control. This was achieved by developing modular, transparent design concepts that were appropriate for the phenomena. The modularity allowed large systems to be designed piecemeal, allowed system failures to be diagnosed and repaired, and allowed multiple stakeholders to participate in the evolution and regulation of systems. We are far from such design concepts in the current stage of development of AI. Moreover, chemical engineering and electrical engineering had at their foundations Schrödinger’s equation and Maxwell’s equations, solid foundations that could guide the development of simplifying modular approximations in the face of exceedingly complex phenomena.
For AI, we certainly have exceedingly complex phenomena—cognitive, social, commercial, and scientific—but we do not have the equivalent of Maxwell’s equations as a guide. We are winging it. Going forward, we therefore need the very best of our overarching, hard-won general scientific and humanistic principles—including rationality, experimentation, dialog, openness, cooperation, skepticism, empathy, and humility—as daily companions on the journey ahead.
Appendix A Inference and Privacy in Databases
We discussed three database design scenarios in Section 2. Figure 4 provides a visual depiction of these scenarios.
In Figure 4(a), we depict the standard privacy-preservation problem, in which the problem is to develop a (randomized) algorithm $Q$ that provides privacy guarantees while ensuring that the privatized response $\tilde{y}$ to a query $x$ is not too far from the “true response” $y$. The latter refers to the response that would be obtained without privacy constraints.
Note that the diagram is not meant to suggest that the randomized algorithm $Q$, which concretely could be a differentially private operator, needs to involve transforming the entire database. Rather, using standard ideas in the differential privacy and online learning literatures (cf. Dwork and Roth, 2014), we envision $Q$ as being called on an as-needed, possibly query-dependent basis.
Figure 4(b) presents an inferential perspective, in which the database is assumed to arise from an underlying population, under a (randomized) sampling operator $S$. The sampling operator may be given or designed (or it may be some blend). The goal is to ensure that $y$ is not too far from the “true response” $y^{*}$. Here the latter is an unobservable population quantity. Statistical theory is needed to make this kind of guarantee.
Figure 4: Database design problems that highlight the complementary nature of computational thinking and inferential thinking.
Of course, in real-world problems there is often a need to combine these two thinking styles. A more general setup is one in which the goal is predictive inference for a new patient, but where we also want to protect the privacy of the individuals in the original dataset, perhaps because their offspring will have an interest in ensuring their own privacy. Figure 4(c) presents the privacy-preserving-inference problem, where we wish to make predictions for new individuals who are sampled from the underlying population while providing a privacy guarantee for the individuals who were in the original database. Here, we want $\tilde{y}$ to be close to $y^{*}$ (see, e.g., Blum et al., 2008; Duchi et al., 2014).
Privacy appears in several of our vignettes, including the database example and the three-layer data market. This is not surprising, given its prominence in public discussion of information technology, including in legal circles, and given that differential privacy and other computational procedures have provided ways to quantify certain aspects of privacy. As we’ve noted, this opens up a discussion of tradeoffs involving economic value and inferential value.
It’s also worth pointing to another computational framework for treating issues surrounding privacy that is known as contextual integrity (Nissenbaum, 2004). Contextual integrity aims to provide formal reasoning strategies for evaluating data flows in social contexts. This is a different direction than the quantitative and algorithmic computation-economics-inference focus of the current paper, but the goals of connecting social desiderata to technological design are shared, as is the emphasis on context. It is of significant interest to engage in a deeper comparison of the current perspective with that of contextual integrity, not only for privacy but for related economic and personal desiderata in AI systems.
Finally, there is much to discuss regarding the interactions of privacy with incentives and trust. This is a domain in which hardware and cryptography can provide some assistance—if an individual is assured that his or her data is being processed using some form of trusted execution environment and/or secure multi-party communication algorithm (Liao et al., 2025), then that individual can be viewed as incentivized to tolerate additional privacy risk.
Appendix B The Missing Middle Kingdom in AI Education
What are the implications of the tripartite blend for undergraduate and graduate education? A naive response would be that students should simply take a collection of classes in computer science, economics, and statistics. But this response is lacking in several ways. First, these three disciplines have not cornered the market on “computational thinking,” “economic thinking” and “inferential thinking.” Important instances of these thinking styles can be found throughout academia. Second, classes in computer science, economics, and statistics often aim to solve problems from previous eras, and the tripartite blend aims at emerging problems. Third, by simply asking students to take more classes we are putting the onus on the students to put together the concepts they are learning in the context of real-world problems. Educators should be shouldering more of the burden of innovative thinking in an era of major change.
It is noteworthy that the phrase “computational thinking” arose in part to support an expanded presence of computational ideas in academic curricula (Wing, 2006). In 2015 my colleagues and I at UC Berkeley built on that perspective to design a new course, Data 8, that aimed to introduce college freshmen to a blending of “computational thinking” and “inferential thinking.” To give a concrete example of what this blend can mean at the freshman level, consider a problem in which the goal is decide whether a treatment is effective based on a comparison of two groups of outcomes. Students were presented with two columns of numbers, labeled “treatment” and “control,” and asked whether it seemed plausible these columns were significantly different (i.e., there was an effect of the treatment), or not. To aid this assessment, students first learned how to make histograms (learning just enough Python programming to carry out this task). Comparing the histograms visually gave a sense of possible differences. To take a step further and address an inferential question, we introduced the permutation test. This test involves some “what-if” thinking, where one considers the possibility that the columns of numbers come from the same source. In that world, one can lump the two columns together, permute the resulting single column, split it into two new columns, and compute whatever comparative statistic between the columns that one likes. (We used the variation distance, which is a simple comparison between histograms). Doing this repeatedly yields a sampling distribution, which can be plotted as a histogram. Placing the statistic of the original two columns on this histogram, one can assess whether this statistic seems likely given the hypothetical “what-if” thinking that the simulation process embodies. Finding that statistic out in the tail of the histogram suggests that that hypothetical assumption may be false.
Clearly there is a lot of inferential reasoning going on in this example, and it can be blended with computational concepts in ways that go beyond just plotting histograms. For example, how does one permute a column of numbers in Python? There is a naive $O(n^{2})$ algorithm to do it, but there is also a more sophisticated $O(n)$ algorithm. Aha! Then, critically, with this blend in hand, we returned to the students and asked them to find their own problems, and accompanying datasets, where “treatment” and “control” were whatever real-world distinction they want to investigate. (E.g., does water quality vary between more wealthy parts of a city and well wealthy parts? or Was there more deforestation in the Amazon this past year as compared to five years ago?). Such real-world investigation was exciting for a diverse set of students, in ways that the problems studied in classical computer science curricula are sometimes not (e.g., game-playing programs). Data 8 became the fastest-growing class in the history of UC Berkeley, and is currently taken by over 1500 students each semester.
Returning to the original design process for Data 8, it is natural to consider adding “economic thinking” to the mix. As I have emphasized in this article, in many problems data arises out of interactions among humans, and the algorithms that analyze such data should take such interactions into account. One can well imagine that basic ideas such as matching markets, contract design, and auctions could be integrated into the algorithms-for-real-world-problem-solving perspective of Data 8.
Such material would retain the core focus on algorithms, but it would also open up coverage of social issues (such as privacy, fairness, ownership, alignment, reputation, and transparency). It would also demonstrate that the world of algorithm design—the key leverage point that we have on the evolution of modern technology—is worth understanding by a broad swath of students, not just computer engineers. Understanding the role of algorithms that address economic and inferential issues can help engender a dialog that is more multi-directional than simply teaching computational concepts to constituencies outside of computer science.
A related virtue of the tripartite blend is that it is inclusive of the wide range of disciplines that have already developed connections to computer science, economics, and statistics. Indeed, from these three starting points, one arrives quickly at disciplines such as cognitive science, social psychology, law, sociology, public policy, mathematics, physics, biology, and the humanities. These are all disciplines that can provide additional perspectives that are essential in shaping human-centric technology. To take but one example where these perspectives link to our discussion, consider behavioral economics, which links cognitive and social psychology with economics, aiming to overcome a perceived overreliance on unrealistic assumptions of rationality in classical economics. Here too, however, existing concepts need to be revisited in the data-centric era. Recalling that humans provide the data on which AI artifacts such as LLMs are based, we see that behavioral data is in some sense already baked into learning-based systems. Blending the massive-but-messy information in such data with the more traditional, experiment-driven insights coming from behavioral economics is an important agenda item for future research.
It is often suggested that there should be stronger links between computer science and the humanities. But in practice this often boils down to having computer science students take an ethics class or another elective, and to having humanities students acquire programming skills. The gap between computer science and the humanities is large, and it is difficult to design direct bridges. This is due in part to the absence of connections that flow through the social sciences in attempts to design such bridges. Indeed, some social sciences have natural connections to computer science (e.g., economics), others have natural connections to the humanities (e.g., history), and in general the focus on humans and groups of humans provides essential perspectives. Our tripartite blend can be viewed as providing a new kind of engineering core that aims to catalyze existing connections and spawn others. It provides a middle kingdom between engineering and the humanities that can help refresh academic debates and help design forward-looking, inclusive curricula.
Appendix C LLMs, Uncertainty, and Collectives
I have argued that the ability to cope with uncertainty—at both the individual level and the collective level—is a hallmark of intelligence. Recall that we are considering a broad, informal notion of uncertainty that involves sampling, information asymmetry, provenance, and other ways in which information that is relevant to a decision may be biased, partially available, or stale.
Current-generation LLMs make little use of any explicit notion of uncertainty. They can give the impression that they are quantifying uncertainty and reasoning inductively, but this is in part an illusion, arising from the fact that the humans who produced the data on which the learning algorithms are based employed terminology and arguments associated with reasoning under uncertainty. It is far from clear that LLMs have extracted useful general principles for managing uncertainty from such data, particularly given that individual humans do not necessarily follow normative rules (see, e.g., Gilovich et al., 2002; Griffiths, 2024). Moreover, some of the known problems with LLMs, such as systematic overconfidence (Sun et al., 2025), are not exhibited by individual humans to the degree that they are exhibited by LLMs. It remains an important item for the research agenda to develop a deeper understanding of the cognitive science of uncertainty and the appropriate decision-theoretic underpinnings for LLMs.
But, as I have argued, a cognitivist perspective is not sufficient for understanding and coping with real-world uncertainty. Moreover, I have suggested that LLMs can be usefully viewed as a collectivist entity rather than a singular intelligence. In this section I provide further discussion of these two issues and their interrelationships.
@@ Appendix C LLMs, Uncertainty, and Collectives
A first point is that collectives play an essential role in changing the world in ways that shape and mitigate uncertainty. An individual human foraging for fruit may or may not succeed in finding fruit on any given day, but once a collective creates a market for produce, the uncertainty regarding finding fruit drops significantly. A human can then depend on fruit being available and build on that certainty, perhaps opening a pastry shop.